OpenStudy (anonymous):

how can I create an equation of the line passing through (3, -1) (3,0) and (3, -3/4) ?

OpenStudy (anonymous):

you first need to pick two points....then you put the two points in to the equation y^2-y^1/x^2-x^1 so it should be 0+1/3-3, then you get your slope which is 1 Next you do the formula y-y1-mx-x1, so y+1=1x-3, then simplify to get y=1x-4

OpenStudy (anonymous):

You understand or still need help??

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

by the way the formula should be y-y1=mx-x1, just a typo

OpenStudy (anonymous):

got it, i forgot the formula. its been a while.